OpenStudy (anonymous):

How are the functions y = x and y = x – 3 related? How are their graphs related?

OpenStudy (anonymous):

The functions are both linear functions and their derivatives are equal Graphically, both are straight lines which are parallel

OpenStudy (anonymous):

in case you aren't doing derivatives yet, @him1618 means that they have the same slope. Notice that the coefficient of x in both functions is 1. a line written as f(x) = mx + b will have slope of m Both these functions have "1" as the x coefficient, so they have the same slope, 1. That makes them parallel. They are separated a bit on the graph... that's because one is y = x - 3, so that y will be equal 0 when x = 3. At the same spot on the x axis, the other line will be at y = 3.

OpenStudy (anonymous):

tell your math teacher (gently) that \(y=x\) is not a function, it is an equation

OpenStudy (anonymous):

they both represent straight lines of the same slope but the line y=x-3 is shifted 3 units along negative y axis

OpenStudy (anonymous):

@satellite73 love it :) "...gently..."

OpenStudy (anonymous):